Calculate the DTI ratio (in percent) for a borrower that has a gross monthly income of $4,775 and has a minimum credit card bill of $115, a car lease payment of $276.78, and a student loan payment of $216, and who is applying for a loan with a monthly payment of $2,021.79. (round your answer to the tenth of the nearest percent).
Answer #1

Gross monthly income = 4775

Monthly debt figure = 115 + 276.78 + 216 + 2021.79 = $ 2,629.57

DTI ratio = (2,629.57/4775) = 55.07%

DTI ratio without the new loan = (115 + 276.78 + 216)/ 4775 = 12.73%